OpenStudy (anonymous):

When 7 is added to three times of a number, the result is 73. Find the number (Let x be the number)

OpenStudy (diyadiya):

Let x be the number 3times of a number =3x when 7 is added it'll become 3x+7 so 3x+7=73

OpenStudy (anonymous):

x=?

OpenStudy (diyadiya):

subtract 7 from both the sides 3x+7-7=73-7 3x=66 divide both sides by 3 3x/3=66/3 x=66/3 x=?

OpenStudy (diyadiya):

What is 66/3 ?

OpenStudy (anonymous):

111111111

OpenStudy (diyadiya):

When you do 66/3 you'll get the value of x

OpenStudy (diyadiya):

@michu

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Yes thank you i love you

OpenStudy (diyadiya):

Sorry its not 11

OpenStudy (anonymous):

????????????????????????????????????????????????????????//

OpenStudy (anonymous):

o no D:

OpenStudy (anonymous):

LIAR

OpenStudy (diyadiya):

Wait ..i'm not lying ..you did a mistake while dividing 66/3

OpenStudy (anonymous):

22

OpenStudy (anonymous):

My math is pellet bro i still love you :D

OpenStudy (diyadiya):

Right! its 22 aand i'm not bro!

OpenStudy (anonymous):

k SIR :D i love you